# Exploring Precision and Recall to assess the quality and diversity of LLMs

### v2

@@ 5.1 Experimental settings
For the models, we stress the distinction between pre-trained models and instruction-tuned (or chat) models. The former are models that have been trained on a large corpus of text, and the latter have been fine-tuned with instructions and optionally aligned with human feedback.
@@
  • Llama-2 {7, 13, 70}B counterparts fine-tuned and aligned with human feedback (Touvron et al., 2023).
@@
  • Mistral 7B and its insctruction-tuned version Mistral 7B Instruct (Jiang et al., 2023).
  • Pythia {6.9,12}B (Biderman et al., 2023)
@@
Figure 7: Correlation between Precision and Recall for Wikipedia’s biographies generation.
@@
We task the models with generating creative texts based on a set of 50 manually-crafted creative instructions such as Write about a dream you had or Write a script for a short film. We leverage the stochastic decoding procedure of models to generate a large number of output based on these 50 prompts. The specificity of this task is that no reference dataset is available.
Appendix B provides full details on experiments.
@@ 5.2 Preliminary results
We first present some preliminary results to illustrate the potential of our metrics.
@@
We consider a reference distribution $P$ made up of journal articles on two topics (World, Sport), extracted from the AG news dataset (Zhang et al., 2015). We then simulate three models. $Q_{1}$ corresponds to a model that generates texts on a single topic. $Q_{2}$ generates texts from the two reference topics. $Q_{3}$ generates text on divergent topics compared to $P$. When considering $P$ as a reference, $Q_{1}$ therefore suffers from a lack of diversity and $Q_{3}$ from a lack of quality. The experiment is illustrated in Figure 2. Precision and Recall metrics allow us to clearly distinguish between these cases, while,for instance, MAUVE would give the same score for both, demonstrating that our metrics can capture the trade-off between quality and diversity on a content-based level.
@@
Figure 7 shows the correlation between Precision and Recall and other distribution-based metrics on the Wikipedia Biography dataset in our experiments. We observe that 1) Recall expectedly reflects word-level diversity as indicated by its correlation with Self-BLEU and Distinct-N 2) Precision shows no correlation with these metrics but does correlate with MAUVE. This suggests that Precision captures the adequacy of distributions, but not diversity, essentially indicating the quality of the generated texts. A complementary analysis of MAUVE on the WebText dataset is available in Appendix B.
@@ 5.3 Evaluation of open-ended generation
Increasing the number of in-context examples make Chat models more diverse.
@@
After a few numbers of in-context examples, we observe that both the Precision and the Recall plateau. This suggests that the model has captured the distribution of the prompt after few examples and that adding more does not bring additional information.
Llama-2 70B Chat Llama-2 13B Chatct Llama-2 7B Chat Mistral 7B Instruct Vicuna 7B
@@
Precision $0.82\pm 0.07$ $0.77\pm 0.08$ $0.78\pm 0.09$ $\mathbf{1.00\pm 0.00}$ $0.67\pm 0.01$
Recall $0.98\pm 0.01$ $0.93\pm 0.01$ $0.91\pm 0.01$ $\mathbf{1.00\pm 0.00}$ $0.95\pm 0.01$
@@
Table 1: Relative comparison of models based on Precision and Recall on the Creative text generation task. Texts generated by the first models are used as a reference against which are compared the others. The values are averaged over 5 generation seeds and the standard deviation is given. The model used as a reference distribution is in bold.
@@ 5.5 Creative text generation
Compared to Llama-2 70B Chat, Mistral 7B Instruct has an almost perfect Precision but suffers from a lower Recall. Geometrically, this means that its distribution fits almost perfectly in the support of Llama-2 70B Chat, but does not cover it totally. Table 1b illustrates even further this phenomena: compared to Mistral 7B Instruct, every model has a quite low precision but a high recall. Geometrically, the Llama-2-based covers a wide portion of the support of Mistral 7B Instruct, but also assigns mass to points outside this distribution.
@@ 6 Conclusion
We introduced Precision and Recall as automated metrics for comparing generative model outputs against a reference distribution. These metrics serve as independent indicators of the quality and diversity of generated text. In challenging scenarios where evaluation protocols are scarce, such as tasks with high generation complexity, we demonstrated the efficacy of our metrics in assessing the quality and diversity of model outputs. Furthermore, our metrics revealed nuanced model behaviors that lie beyond the scope of traditional evaluation benchmarks. We believe our work marks a significant step towards advancing the evaluation of generative models, offering insights to better understand their capabilities and facilitating the development of more sophisticated evaluation protocols within the community.
@@ 7 Limitations
Our study proposes to adapt the framework of Kynkäänniemi et al. (2019) from image to text generation. However, limitations of the orinal image-based frame work have been highlighted (Naeem et al., 2020; Kim et al., 2023) and may also apply to our adaptation in NLP. Notably, the sensitivity to outliers and the generalizability of these metrics across different domains warrant further investigation.
The metrics proposed in our study rely on an auxiliary model, specifically GPT-2 embeddings, to rate the quality of text generations. While we follow previous work demonstrating successful use of such embeddings, the choice of embedding may significantly impact the performance and applicability of these metrics. Future research could explore the effects of alternative embedding models on the Precision and Recall measurements in text generation tasks.
We present these metrics as valuable new instruments for evaluating a model’s open-ended generation capabilities. Nonetheless, to comprehensively assess LLMs across specific dimensions and tasks, these metrics may be supplemented by other approaches.
@@ 8 Ethical considerations
Precision and Recall are distributional-based metrics, and therefore depend heavily on the data they are tested against. If these data have biases, our metrics might unintentionally favor models that repeat these biases. This issue is critical because it can lead to unfair outcomes, especially for underrepresented groups. This highlights the need for a carefully crafted reference dataset.
To tackle this problem, we suggest a practical solution: calculate Precision and Recall for specific groups within the data, such as different genders or ethnic minorities. This helps to check whether the model is as accurate and fair for one group as it is for another. Our goal is to make sure that the models we are evaluating do not overlook or misrepresent any group. Further investigations should be conducted on this consideration.
@@ Appendix A Definition of Precision-Recall Curves
In this section, we introduce another object called the PR-Curve, which is closely connected to the MAUVE Score. Then, we demonstrate using a straightforward example that the MAUVE score struggles to distinguish between a lack of diversity and a lack of quality. This difficulty can actually be explained by the shortcomings of the underlying curve used by the MAUVE score.
@@ A.1 Definition
Witnessing the need to distinguish between quality and diversity, and making the parallel with the notion of precision and recall from classification, Sajjadi et al. (2018) introduced a new definition of Precision and Recall for discrete distributions.
@@
A distribution $Q$ has a precision $\alpha$ and a recall $\beta$ with respect to a distribution $P$ if there exists the distributions $\mu$, $\nu_{P}$ and $\nu_{Q}$ such that: $P=\alpha\mu$ $+(1-\alpha)\nu_{P}$ (3a) $Q=\beta\mu$ $+(1-\beta)\nu_{Q}.$ (3b)
@@
This definition is illustrated in Figure A.1. The distribution $\nu_{P}$ represents the part of $P$ that cannot be generated by $\mu$ and thus $Q$, and $\nu_{Q}$ represents the part of $P$ that should not be generated by $\mu$. For every distribution $\mu$ Figure A.1 defines a couple, i.e. a trade-off, $(\alpha,\beta)$, and we can define the Precision Recall Curve as the set of all Pareto-optimal trade-offs. More recently, Simon et al. (2019) has extended the definition for any continuous distribution $P$ and $Q$:
@@
Intuitively, for any given trade-off parameter $\lambda\in[0,+\infty]$, the precision $\alpha_{\lambda}$ diminishes when $q({\bm{x}})$ is smaller than $\lambda p({\bm{x}})$ at certain points ${\bm{x}}$, indicating overestimation of regions by the distribution $Q$. On the contrary, the recall $\beta_{\lambda}$ declines when $p({\bm{x}})$ is smaller than $\lambda q({\bm{x}})$ for specific points ${\bm{x}}$, suggesting underestimation of regions by $Q$. Although the PR-Curve is valuable for evaluating the quality and diversity of generative models, its interpretation can be challenging.
More concretely, an example of PR-Curves is depicted in Figure A.2. Figure A.2a shows samples drawn from the reference distribution $P$. Here, $Q_{1}$ represents a distribution capable of generating high-quality images, albeit only for the first 5 digits, while $Q_{2}$ represents a distribution that generates noisy images on all labels. The PR-Curve of $Q_{1}$ surpasses that of $Q_{2}$ for high values of $\lambda$ (top left), indicating the superior quality of $Q_{1}$. Conversely, for low values of $\lambda$ (lower right), the PR-Curve of $Q_{2}$ exceeds that of $Q_{1}$, indicating greater diversity in $Q_{2}$.
For a two-number summary of these curves, Sajjadi et al. (2018) has proposed the $F_{\gamma}$-scores: $\mathrm{F}_{\gamma}=\max_{\lambda}(1+\gamma)^{2}\alpha_{\lambda}\beta_{\lambda}/(\gamma^{2}\alpha_{\lambda}+\beta_{\lambda})$. The $\mathrm{F}_{1/8}$ is employed to assess the quality of the model, and $\mathrm{F}_{8}$ score is used to assess the diversity. Abusing the notation, the scores $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ are often refered as the Precision and the Recall. Note that in this work, we use another two-number summary of the PR-Curve - $\alpha_{\infty}=Q(\mathrm{Supp}(P))$ and $\beta_{0}=P(\mathrm{Supp}(Q))$ which is practically easier to estimate.
@@
Figure A.2: PR-Curves for the MNIST dataset. The target distribution $P$ is shown in (a), and the distributions $Q_{1}$ and $Q_{2}$ are shown in (b) and (c) respectively. The PR-Curves of $Q_{1}$ and $Q_{2}$ are shown in (d). The PR-Curve of $Q_{1}$ indicates high quality and limited diversity and the PR-Curve of $Q_{2}$ indicates limited quality but high diversity.
A.2 Connection with the MAUVE score
The PR-Curve bears a close relationship with the MAUVE score Pillutla et al. (2021, 2023), which represents the area under the curve of a modified version of the PR-Curve introduced by Djolonga et al. (2020). On our AG News experiment, fully described in Section 5.2, we can exemplify some limitations of this approach. Specifically, Figure A.3 shows that the underlying curves used for computing the MAUVE score can hardly differentiate the discrepancy highlighted in Figure 2, when the original PR-Curves from Sajjadi et al. (2018) can. When adapting the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores to the MAUVE curves, and plotting the results on Figure A.4 we find that these adapted scores are substantially less meaningful than those from (Sajjadi et al., 2018).
Figure A.3: Comparison of the PR-Curve and the MAUVE score for the AG News dataset. The PR-Curve is a finer representation of the MAUVE score. However, the PR-Curves introduced by Sajjadi et al. (2018) clearly reflects the lack of diversity or lack of quality.

### v3

@@ 5.1 Experimental settings
For the models, we stress the distinction between pre-trained models and instruction-tuned (or chat) models. The former are models that have been trained on a large corpus of text, and the latter have been fine-tuned with instructions and optionally aligned with human feedback. More details on the experimental settings are provided in Appendix B.
@@
  • Llama-2 {7, 13, 70}B and their counterparts fined-tuned and aligned with human-feedback (Touvron et al., 2023).
@@
  • Mistral 7B and its instruction-tuned version Mistral 7B Instruct (Jiang et al., 2023).
  • Pythia {6.9,12}B (Biderman et al., 2023).
@@
(a) Correlation between Precision, Recall, MAUVE and traditional metrics.
(b) Correlation between Precision, Recall, MAUVE and the ChatGPT-3.5-turbo metrics.
Figure 7: Correlation with Precision, Recall and MAUVE with different metrics for Wikipedia’s biographies generation. Precision is on average positively correlated to quality evaluation, while Recall is positively correlated with diversity evaluation.
@@
We task the models with generating creative texts based on a set of 50 manually-crafted creative instructions such as Write about a dream you had or Write a script for a short film. We leverage the stochastic decoding procedure of models to generate a large number of outputs based on these 50 prompts. The specificity of this task is that there is no reference dataset available.
@@ 5.2 Preliminary results
First, we present some preliminary results to illustrate the potential of our metrics.
@@
We consider a reference distribution $P$ made up of journal articles on two topics (World, Sport), extracted from the AG news dataset (Zhang et al., 2015). We then simulate three models. $Q_{1}$ corresponds to a model that generates texts only from the World topic. $Q_{2}$ generates texts from the two reference topics. $Q_{3}$ generates text on divergent topics compared to $P$. When considering $P$ as a reference, $Q_{1}$ therefore suffers from a lack of diversity and $Q_{3}$ from a lack of quality. The experiment is illustrated in Figure 2. Precision and Recall metrics allow us to clearly distinguish between these cases, while, for instance, MAUVE would give the same score for both, demonstrating that our metrics can capture the trade-off between quality and diversity on a content-based level.
@@
Figure 7a shows the correlation between Precision and Recall and other distribution-based metrics on the Wikipedia Biography dataset in our experiments. We observe that:
  1. 1.
    Recall expectedly reflects word-level diversity, as indicated by its correlation with Self-BLEU and Distinct-4.
  2. 2.
    Precision shows no correlation with these metrics, but does correlate with MAUVE.
This suggests that Precision captures the adequacy of distributions, but not diversity, essentially indicating the quality of the generated texts. A complementary analysis of MAUVE on the WebText dataset is available in Section B.4.
@@ 5.3 Evaluation of open-ended generation
Increasing the number of in-context examples make "chat" models more diverse.
@@
After a few numbers of in-context examples, we observe that both the Precision and the Recall plateau. This suggests that the model has captured the distribution of the reference datasets after a few examples and that adding more examples does not bring additional information.
@@
Mistral 7B Instruct Llama-2 70B Chat Llama-2 13B Chat Llama-2 7B Chat Vicuna 7B
Precision $\mathbf{1.00\pm 0.00}$ $0.82\pm 0.07$ $0.77\pm 0.08$ $0.78\pm 0.09$ $0.67\pm 0.01$
Recall $\mathbf{1.00\pm 0.00}$ $0.98\pm 0.01$ $0.93\pm 0.01$ $0.91\pm 0.01$ $0.95\pm 0.01$
@@
Table 1: Relative comparison of models based on Precision and Recall on the Creative text generation task. The texts generated by the first models are used as a reference against which are compared the others. The values are averaged over 5 generation seeds and the standard deviation is given. The model used as a reference distribution is in bold.
@@ 5.5 Creative text generation
Compared to Llama-2 70B Chat, Mistral 7B Instruct has an almost perfect Precision but suffers from a lower Recall. Geometrically, this means that its distribution fits almost perfectly in the support of Llama-2 70B Chat, but does not cover it totally. Table 1b illustrates even further this phenomena: compared to Mistral 7B Instruct, every model has a quite low precision but a high recall. Geometrically, the Llama-2-based models cover a wide portion of the support of Mistral 7B Instruct, but also assigns mass to points outside this distribution.
5.6 Correlation with human-like preferences
Finally, we evaluated how our proposed metrics relate to human evaluation. Since one of our evaluation involve significantly long inputs, we conducted a mock-human evaluation using GPT-3.5-turbo. To evaluate the quality, we ask GPT-3.5-turbo to rate our generated biographies on three dimensions: coherence, insightfulness and Wikipedia-likeness. We also experimented with a proxy of diversity, by asking GPT-3.5-turbo to rate the similarity of batches of 10 texts (i.e., low score if there are many similar texts). More details regarding the experimental setup are given in Section B.2.
Precision Correlates with GPT-3.5 quality evaluation.
Models with high precision generate texts that align closely with the reference distribution, which consists of high-quality Wikipedia articles. Consequently, higher precision typically corresponds to better GPT-3.5 scores, as illustrated in Figure 8. Note that instruction-tuned models, which are tuned similarly to GPT-3.5, were excluded from the correlation analysis with GPT-3.5 quality scores due to their uniformly high scores, relatively to pre-trained models.
Recall correlates with human-like diversity evaluation.
As depicted on Figure 8, we found a positive correlation between our introduced Recall and human-like evaluation of diversity. This confirms our findings regarding the generated contents of LLMs.
@@ 6 Conclusion
We introduced Precision and Recall as automated metrics for comparing generative model outputs against a reference distribution. These metrics serve as independent indicators of the quality and diversity of the generated text. In challenging scenarios where evaluation protocols are scarce, such as tasks with high generation complexity, we demonstrate the efficacy of our metrics in assessing the quality and diversity of model output. Furthermore, our metrics revealed nuanced model behaviors that lie beyond the scope of traditional evaluation benchmarks. We believe that our work contributes to advancing the evaluation of generative models, offering valuable insights to better understand their capabilities, and aiding in the development of more refined evaluation protocols within the community.
@@ 7 Limitations
Our study proposes to adapt the framework of Kynkäänniemi et al. (2019) from image to text generation. However, limitations of the original image-based framework, as highlighted by Naeem et al. (2020) and Kim et al. (2023), may also apply to our adaptation to NLP. In particular, the sensitivity to outliers and the generalizability of these metrics across different domains warrant further investigation.
The metrics proposed in our study are based on an auxiliary model, specifically GPT-2LARGE embeddings, to rate the quality of text generations. Although the metrics tend to be correlated across models as shown in Section B.1, the choice of embeddings may still significantly impact the performance and applicability of these metrics. Future research could explore the effects of alternative embedding models on the Precision and Recall measurements in text generation tasks.
We present these metrics as valuable tools for evaluating a model’s open-ended generation capabilities. However, to comprehensively assess LLMs across specific dimensions and tasks, these metrics should be supplemented by other approaches. In particular, the relationship with traditional human evaluation needs to be investigated more thoroughly in future research. Although we performed a mock human evaluation using GPT-3.5-turbo, this is not a complete assessment of how these metrics compare to the human evaluation.
@@ 8 Ethical considerations
Precision and Recall are metrics that rely heavily on the distribution of the data they are tested against. If the data contain biases, these metrics may favor models that perpetuate those biases. This issue is crucial because it can lead to unfair outcomes, particularly for underrepresented groups, emphasizing the need for a well-designed reference dataset.
To address this problem, we propose a practical solution: calculate Precision and Recall for specific groups within the data, such as different genders or ethnic minorities. This approach helps to ensure that the model is equally accurate and fair between all groups. Our objective is to prevent models from overlooking or misrepresenting any group. Further investigations should focus on this consideration.
@@ Appendix A Definition of Precision-Recall Curves
In this section, we introduce a mathematical tool called the PR-Curve, which is used to evaluate generative models. PR-Curves are closely related to the MAUVE score. We will then list the main differences between PR-Curves and the MAUVE score to highlight the limitations of the latter. Following this, we will demonstrate with a straightforward example that the MAUVE score struggles to distinguish between a lack of diversity and a lack of quality. This difficulty can be explained by some of the distinguishing points between PR-Curves and the MAUVE score.
@@ A.1 Definition
Recognizing the need to distinguish between quality and diversity, and drawing a parallel with the concepts of precision and recall from classification, Sajjadi et al. (2018) introduced a new definition of Precision and Recall for discrete distributions.
@@
A distribution $Q$ has a precision $\alpha$ and a recall $\beta$ with respect to a distribution $P$ if there exists distributions $\mu$, $\nu_{P}$ and $\nu_{Q}$ such that:
$P=\alpha\mu$ $+(1-\alpha)\nu_{P}$ (3a)
$Q=\beta\mu$ $+(1-\beta)\nu_{Q}.$ (3b)
@@
This definition is illustrated in Figure A.1. The distribution $\nu_{P}$ represents the part of $P$ that cannot be generated by $\mu$ and thus $Q$, and $\nu_{Q}$ represents the part of $Q$ that should not be generated by $\mu$. The set of all attainable couples $(\alpha,\beta)$ such that there is a distribution $\mu$ from Equation 3 defines various trade-offs. The Precision Recall Curve can then be defined as the set of all Pareto-optimal trade-offs. More recently, Simon et al. (2019) has extended the definition for any continuous distribution $P$ and $Q$:
@@
Intuitively, for any given trade-off parameter $\lambda\in[0,+\infty]$, the precision $\alpha_{\lambda}$ diminishes when $q({\bm{x}})$ is smaller than $\lambda p({\bm{x}})$ at certain points ${\bm{x}}$, indicating overestimation of regions by the distribution $Q$. On the contrary, the recall $\beta_{\lambda}$ declines when $p({\bm{x}})$ is smaller than $\lambda q({\bm{x}})$ for specific points ${\bm{x}}$, suggesting underestimation of regions by $Q$. More concretely, an example of PR-Curves is depicted in Figure A.2. Figure A.2a shows samples drawn from the reference distribution $P$. Here, $Q_{1}$ represents a distribution capable of generating high-quality images, albeit only for the first 5 digits, while $Q_{2}$ represents a distribution that generates noisy images on all labels. The PR-Curve of $Q_{1}$ surpasses that of $Q_{2}$ for high values of $\lambda$ (top left), indicating the superior quality of $Q_{1}$. In contrast, for low values of $\lambda$ (lower right), the PR-Curve of $Q_{2}$ exceeds that of $Q_{1}$, indicating greater diversity in $Q_{2}$.
For a two-number summary of these curves, Sajjadi et al. (2018) has proposed the $F_{\gamma}$-scores: $\mathrm{F}_{\gamma}=\max_{\lambda}(1+\gamma)^{2}\alpha_{\lambda}\beta_{\lambda}/(\gamma^{2}\alpha_{\lambda}+\beta_{\lambda})$. The $\mathrm{F}_{1/8}$ is employed to assess the quality of the model, and $\mathrm{F}_{8}$ score is used to assess diversity. Abusing the notation, the scores $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ are often referred as Precision and Recall. In practice, the PR-Curves are estimated by sampling from the distributions $P$ and $Q$, and by performing density estimation on the embeddings on the samples. There are several ways to estimate the PR-Curves, including a $k$-means estimator (Sajjadi et al., 2018) or a discriminator method (Simon et al., 2019).
Connection with PR-Curves and Precision/Recall:
In this paper, we have used another two-number summary of the PR-Curve introduced by Kynkäänniemi et al. (2019): the maximum precision $\alpha_{\infty}$ and the maximum recall $\beta_{0}$. As a matter of fact, one can show that:
$\alpha_{\infty}=\max_{\lambda}\alpha_{\lambda}=Q(\mathrm{Supp}(P))\quad\text{and}\quad\beta_{0}=\max_{\lambda}\beta_{\lambda}=P(\mathrm{Supp}(Q)).$
These two values encapsulate the definition of support provided by Kynkäänniemi et al. (2019). The first value, $\beta_{0}=P(\mathrm{Supp}(Q))$, measures the proportion of $P$ covered by the support of $Q$. The second value, $\alpha_{\infty}=Q(\mathrm{Supp}(P))$, measures the proportion of $Q$ covered by the support of $P$. The primary practical advantage is that these two values can be computed without performing a density estimation. As detailed in Section 3, we use a $k$-nearest neighbors estimator to estimate these values.
@@
Figure A.2: PR-Curves for the MNIST dataset. The target distribution $P$ is shown in (a), and the distributions $Q_{1}$ and $Q_{2}$ are shown in (b) and (c) respectively. The PR-Curves of $Q_{1}$ and $Q_{2}$ are shown in (d). The PR-Curve of $Q_{1}$ indicates high quality and low diversity and the PR-Curve of $Q_{2}$ indicates low quality but high diversity.
Figure A.3: Comparison of the PR-Curve from (Sajjadi et al., 2018) (left) and the divergence frontier used to compute MAUVE score (Pillutla et al., 2021) for the AG News dataset. Visually, MAUVE’s curve struggle to distinguish between a lack of diversity and a lack of quality, while the PR-Curves on the left provide a finer evaluation.
A.2 MAUVE score
MAUVE score definition.
The MAUVE score (Pillutla et al., 2021) is a measure of discrepancy between two distributions $P$ and $Q$ in the embedding space. The MAUVE score is defined as the area under the divergence frontier between $P$ and $Q$. The divergence frontier is defined as the set of points $(\alpha_{\pi},\beta_{\pi})$ for $\pi\in[0,1]$:
$\begin{cases}\alpha_{\pi}=\exp\left(-D_{\mathrm{KL}}(Q\|\pi P+(1-\pi)Q)\right)\\ \beta_{\pi}=\exp\left(-D_{\mathrm{KL}}(P\|\pi P+(1-\pi)Q)\right).\end{cases}$
In other words, the MAUVE score is a one-number summary of another quality-diversity curve.
Connection between MAUVE and PR-Curves.
The MAUVE score is based on a curve different from PR-Curves, and there is no straightforward connection between the two. But, from a practical point of view, both curves are estimated using density estimation on the embeddings of samples from $P$ and $Q$. Pillutla et al. (2021) used a $k$-means estimator to estimate the divergence frontier and then computed the area under the curve to obtain the MAUVE score.
Since both curves are based on density estimation, we can compare them using the same dataset, embeddings, and density estimation method. This allows us to highlight the limitations of the divergence frontier used in the MAUVE score. In our AG News experiment, fully described in Section 5.2, we exemplify these limitations. Specifically, Figure A.3 shows that the underlying curves used for computing the MAUVE score struggle to differentiate the discrepancies highlighted in Figure 2, while the original PR-Curves from Sajjadi et al. (2018) can.
When adapting the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores to the MAUVE curves and plotting the results in Figure A.4, we find that these adapted scores are substantially less meaningful than those from Sajjadi et al. (2018). Furthermore, the definitions of Precision and Recall we used, based on support-based metrics (Kynkäänniemi et al., 2019), appear to be more informative than the $\mathrm{F}_{1/8}$ and $\mathrm{F}_{8}$ scores.